In a magnetic field of 14x10^-4 T a 4.2x10^-6 C flows at a velocity of 4x10^6 m/s. The angle is


90 degrees. How strong is the magnetic force operating on the charge?

1
Expert's answer
2022-05-12T15:09:17-0400
"F=qvB\sin\theta\\
=4.2*10^{-6}*4*10^6*14*10^{-4}*\sin90^\circ\\
=0.024\:\rm N"